### v1

# A Hierarchy of Topological and Superconducting States in Rhombohedral Hexalayer Graphene

thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.
A Hierarchy of Topological and Superconducting States in Rhombohedral Hexalayer Graphene
Ron Q. Nguyen1    Hai-Tian Wu1    Erin Morissette1    Naiyuan J. Zhang1    Peiyu Qin1    K. Watanabe2    T. Taniguchi3    Aaron W. Hui1    D. E. Feldman1,5    J.I.A. Li1,4 jia.li@austin.utexas.edu 1Department of Physics, Brown University, Providence, RI 02912, USA 2Research Center for Functional Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 3International Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 4Department of Physics, University of Texas at Austin, Austin, TX 78712, USA 5Brown Theoretical Physics Center, Brown University, Providence, Rhode Island 02912, USA
(July 29, 2025)
Superconductivity and the quantum Hall effect are conventionally viewed as mutually exclusive: the former is suppressed by magnetic fields, while the latter relies on them. Here, we report the surprising coexistence of these two phenomena in rhombohedral hexalayer graphene. In this system, a superconducting phase is not destroyed—but instead stabilized—by an out-of-plane magnetic field. Strikingly, this superconducting state coexists and competes with a sequence of quantum Hall states that appear at both integer and half-integer Landau level fillings. Both the superconducting and quantum Hall states exhibit sharply defined thermal transitions or crossovers, with nearly identical onset temperatures—pointing to a shared underlying mechanism. Taken together, our observations uncover an unprecedented interplay between superconducting and topological phases, challenging conventional paradigms and opening a new frontier in condensed matter physics.
Figure 1: Multi-dimensional phase space of rhombohedral hexalayer graphene. (a) Schematic of the R6G heterostructure. The red arrow indicates the out-of-plane magnetic field $B_{\perp}$. (b) Schematic phase diagram defined by carrier density $n$, displacement field $D$, and out-of-plane magnetic field $B_{\perp}$. (c, d) $n$$D$ maps of longitudinal resistance measured at $B=0$, with current flowing along the principal axes of (c) minimum and (d) maximum conductivity. (e) Annotated $n$$D$ map highlighting prominent phases corresponding to panels (c, d). (f, g) $n$$B$ maps measured at $D=967$ mV/nm, showing (f) longitudinal resistance $R_{xx}$ and (g) transverse resistance $R_{xy}$. White dashed lines trace the trajectories of integer quantum Hall states, exhibiting integer Streda slopes. The boundaries of regime I, highlighted by the black dashed lines, follow a Streda slope of $t=2.5$. In panel (f), white dotted lines show Streda slopes of 2 and 3, which clearly deviate from the boundaries of regime I.
Figure 2: $B_{\perp}$-stabilized superconductivity and coexisting quantum Hall state. $n$$D$ maps of (a) $R_{xx}$ and (b) $R_{xy}$ measured at different values of $B_{\perp}$. The dashed line outlines the boundary of regime I, while the dash-dotted line encloses the superconducting phase. (c) Line cuts of $|R_{xy}|$ as a function of $n$ at $D=1000$ mV/nm, measured at various values of $B_{\perp}$. (d, e) Temperature dependence of $R_{xx}$ (red) and $R_{xy}$ (blue) measured from (d) the $B_{\perp}$-stabilized superconducting phase and (e) the adjacent quantum Hall state.
Figure 3: A sequence of half-integer quantum Hall states. (a, b) $n$$B$ map of $R_{xx}$ (top panel) and $R_{xy}$ (bottom panel) measured at (a) $D=980$ mV/nm and (b) $D=1000$ mV/nm. The superconducting region is enclosed by a white dashed line in the top panel, while the sequence of quantum Hall states forms a Landau fan, marked by black dashed lines. (c, d) Temperature dependence of $R_{xy}$ measured at different $B_{\perp}$ values within (c) the quantum Hall states at both integer and half-integer Landau level fillings and (d) the superconducting state.
[…]
When superconductivity persists to high magnetic fields, it opens the door to a rare and intriguing regime in which superconductivity coexists with another hallmark of two-dimensional (2D) electron systems: the quantum Hall effect. The interplay between these two phenomena has long been proposed as a key platform for realizing fault-tolerant quantum computation, owing to the potential emergence of non-Abelian anyons and topological superconductivity Nayak et al. (2008); Das Sarma et al. (2015); Peralta Gavensky et al. (2020); Alicea and Fendley (2016). Over the past decade, this coexistence has been explored extensively in engineered hybrid systems, where superconducting and quantum Hall regions are coupled across artificial interfaces Amet et al. (2016); Lee et al. (2017); Gül et al. (2022); Zhao et al. (2023), as well as in 2D material heterostructures hosting superconductivity and topological order at zero magnetic field Stepanov et al. (2020); Choi et al. (2025). Despite these advances, achieving an intrinsic intertwinement of superconductivity and the quantum Hall effect within a single 2D electronic system—without relying on interface engineering—remains a central open challenge. A recent experimental report has identified such coexistence near zero magnetic field, sparking renewed interest in the possibility of anyonic superconductivity and its topological implications  Xu et al. (2025); Nosov et al. (2025); Pichler et al. (2025).
Here, we report an unprecedented form of interplay between superconductivity and topological order under a large out-of-plane magnetic field, realized in a rhombohedral hexalayer graphene (R6G) heterostructure  Zhou et al. (2021a, b, 2022); Han et al. (2025); Choi et al. (2025). In stark contrast to all previously reported superconducting phases in 2D material heterostructures, the superconducting state in R6G is stabilized—rather than suppressed—by an external magnetic field. Strikingly, quantum Hall states emerge in direct proximity to this field-induced superconducting phase and form an unconventional sequence that includes both integer and half-integer Landau level (LL) fillings. Through multiple lines of experimental evidence, we identify a common microscopic origin underlying both phases. These findings place strong constraints on viable theoretical interpretations and point to a highly unconventional physical regime, where superconductivity and quantum Hall order emerge from a unified framework.
Facilitated by a dual-encapsulated heterostructure, as illustrated in Fig. 1a, the low-temperature phase space of R6G is governed by three independently tunable parameters: carrier density $n$, displacement field $D$, and out-of-plane magnetic field $B_{\perp}$. Superconductivity and the sequence of quantum Hall states are both confined within a unique boundary, which encircles a tilted cylindrical volume within this $n$$D$$B$ space, as illustrated in Fig. 1b. For simplicity, we will refer to this portion of the phase space as regime I.
We begin by defining the boundary of regime I through its distinct transport signatures. To this end, we perform measurements across multiple two-dimensional slices of the $n$$D$$B$ parameter space. In particular, we examine $n$$D$ planes at five different values of out-of-plane magnetic field $B_{\perp}$ (blue dashed lines in Fig. 1b), as well as $n$$B$ planes at three fixed values of displacement field $D$ (red dashed lines in Fig. 1b).
At zero magnetic field ($B_{\perp}=0$), the cross section of regime I is bounded by the electronic smectic phase and the striped superconducting state, both previously identified through angle-resolved transport studies Morissette et al. (2025a). The transport response across this regime is characterized in Figs. 1c–d, which show longitudinal resistance measured across the $n$$D$ plane at $B_{\perp}=0$, along two orthogonal principal axes corresponding to maximum and minimum conductivity. The smectic phase occupies the region of pronounced transport anisotropy (see Fig. M2b), highlighted by the red-shaded area in Fig. 1e. Within this anisotropic regime, two superconducting phases—denoted SC i and SC ii—emerge concurrently with the smectic order and inherit its strong directional character  Morissette et al. (2025a).
The $B_{\perp}=0$ cross-section of regime I appears as a metallic phase situated between SC i and SC ii, enclosed by black dashed lines in Figs. 1c–e. As $B_{\perp}$ increases, the boundaries of regime I trace a well-defined trajectory in the $n$$B$ plane, marked by peaks in the longitudinal resistance $R_{xx}$ and abrupt jumps in the Hall resistance $R_{xy}$. These boundaries, indicated by black dashed lines in Fig. 1f-g, exhibit a Streda slope of $t=2.5$  Středa and Smrčka (1983); Widom (1983); MacDonald and Středa (1984).
Within the stripe regime, two additional metallic phases—regime II and regime III—are highlighted in Figs. 1c–e. These phases resemble regime I in their transport characteristics and exhibit well-defined boundaries in the $n$$B$ plane, following distinct Streda slopes: $t=2$ for regime II and $t=3.5$ for regime III (see Fig. M6 for further discussion). The presence of such sharply delineated, evolving boundaries is intriguing. Although a well-defined Streda slope often signals a topological phase, we do not observe transport features typically associated with incompressible quantum Hall states Huang et al. (2025). We revisit this issue later in the context of additional experimental evidence.
In addition to their distinct Streda slopes, regimes I, II, and III exhibit significantly reduced transport anisotropy compared to the surrounding stripe phase (see Fig. M2c and Fig. M3a), suggesting the emergence of a competing order that disrupts the underlying smectic state.
Figures 2a–b show $n$$D$ maps of the longitudinal and Hall resistances measured at various values of $B_{\perp}$. At each field, regime I is delineated by peaks in $R_{xx}$, defining a boundary that shifts systematically with $B_{\perp}$ in both carrier density $n$ and displacement field $D$. These evolving boundaries are marked by dashed lines in Figs. 2a–b. Most notably, upon applying a non-zero $B_{\perp}$, a superconducting pocket emerges on the low-density side of regime I, while a sequence of quantum Hall states appears on the high-density side.
The superconducting phase—enclosed by dash-dotted lines—is identified by vanishing longitudinal resistance $R_{xx}$ and zero Hall resistance $R_{xy}$, appearing as dark blue and white, respectively, in the color maps. Figure 2c presents line cuts of the transport response as a function of $n$, taken at fixed $D=1000$ mV/nm and varying $B_{\perp}$. At $B_{\perp}=0$, the system remains resistive across the entire density range. In contrast, at $B_{\perp}=2.5$ and 3.33.33.3 T, a superconducting dome is clearly visible, characterized by a sharp drop to zero resistance. Upon further increasing $B_{\perp}$ to 555 T, the superconducting phase is fully suppressed.
On the high-density side of the superconducting region, we observe the emergence of a distinct phase along the upper boundary of regime I, exhibiting transport signatures that sharply contrast with those of the superconducting state. In this regime, vanishing $R_{xx}$ is accompanied by a pronounced peak in $R_{xy}$. The corresponding conductivity matrix is characterized by vanishing diagonal (longitudinal) components and finite, antisymmetric off-diagonal (Hall) components. In Fig. 3, we show that this high-density state evolves with varying $B_{\perp}$ along a well-defined Streda slope—a hallmark of an incompressible quantum Hall state. Notably, the minima in $R_{xx}$ for each quantum Hall state are flanked by resistive peaks, suggesting that these states arise from strong electronic interactions.
Remarkably, both the superconducting and quantum Hall phases exhibit sharply defined onsets as a function of temperature (Fig. 2d–e). The $R$$T$ curves in Fig. 2d display behavior consistent with a superconducting transition, characterized by a critical temperature $T_{c}\approx 0.5$ K. In contrast, the similarly sharp onset observed for the quantum Hall state in Fig. 2e is highly unusual. Conventional quantum Hall states typically display a gradual thermal activation in $R_{xx}$, while $R_{xy}$ remains largely quantized and temperature-independent (see Fig. M5). The abrupt temperature dependence observed here suggests an unconventional origin for the quantum Hall state, a possibility we return to in light of additional experimental evidence discussed below.
Next, we investigate the unusual robustness of the superconducting phase against $B_{\perp}$ in Figs. 3a–b, which present transport measurements across $n$$B$ planes intersecting regime I at fixed displacement fields. A pronounced resistive peak in $R_{xx}$ appears within regime I around $B_{\perp}=1$ T, marking the lower-field boundary for both the superconducting and quantum Hall states. We interpret this feature as indicative of a transition in the underlying electronic order.
Given that the superconducting phase is stabilized above this resistive feature, it is natural to consider a transition in the underlying isospin order. For $B_{\perp}>1$ T, regime I is expected to be fully polarized in both spin and valley degrees of freedom. In this high-field regime, superconductivity likely originates from Cooper pairing in an odd orbital angular momentum channel, consistent with a $p$-wave pairing symmetry Anderson and Morel (1961). At lower fields ($B_{\perp}<1$ T), the ground state may be only partially polarized. Fermiology analysis reveals that regime I lies between a half-metallic phase at low density and a quarter-metallic phase at high density, consistent with partial isospin polarization in the intermediate regime. However, this interpretation may be at odds with the chiral nature of the SC i and SC ii phases observed near zero field Han et al. (2025), which have been argued to emerge from a fully isospin-polarized normal state.
An alternative interpretation is that the resistive feature marks a $B_{\perp}$-driven reconstruction of momentum-space order. In rhombohedral graphene (R6G), rotational symmetry breaking has been attributed to spontaneous polarization in momentum space Morissette et al. (2025b, a); Dong et al. (2023); Jung et al. (2015); Huang et al. (2023). The stability of the superconducting phase may therefore depend sensitively on the momentum-space configuration of the Fermi surface, particularly within a quarter-metallic state  Geier et al. (2024); Parra-Martinez et al. (2025). This possibility is further supported by the observed $B_{\perp}$-dependent rotation of the principal axis within regime I (see Fig. M3).
On the high-density side of regime I, quantum Hall states emerge along well-defined linear trajectories in the $n$$B$ plane, marked by black dashed lines in Fig. 3a–b. These trajectories extrapolate back to the low-density boundary of regime I at $B_{\perp}=0$, forming a characteristic pattern reminiscent of a Landau fan diagram Zhang et al. (2005). Figure M4a presents a line trace cutting through this sequence of quantum Hall states. The evolution of $R_{xx}$ as a function of both $B_{\perp}$ and $n$ shows excellent agreement with Shubnikov–de Haas (SdH) oscillations, providing further confirmation of a robust quantum Hall sequence.
The slopes of the quantum Hall trajectories allow us to extract the Streda slopes, revealing an intriguing hierarchical structure: the sequence of quantum Hall states occupy both integer and half-integer Landau level (LL) fillings, $\nu=N/2$ with $N\in\mathbb{Z}$.
In conventional two-dimensional electron systems, incompressible states at half-filled LLs—commonly referred to as even-denominator fractional quantum Hall effects (FQHE)—have attracted significant interest due to their non-trivial topological nature. These states typically exhibit much smaller energy gaps than neighboring integer quantum Hall (IQHE) states, and their stability is strongly dependent on $B_{\perp}$.
Interestingly, the quantum Hall states within regime I do not exhibit the expected dependence on $B_{\perp}$. Instead, they show sharply defined temperature onsets that remain nearly constant over a broad range of $B_{\perp}$, as shown in Fig. 3c, indicating that their stability is largely independent of magnetic field. Moreover, the onset temperatures at both integer and half-integer filling factors are nearly identical, suggesting comparable robustness. These observations point to an unconventional hierarchical structure, which is further supported by the Shubnikov–de Haas (SdH) oscillations measured across the quantum Hall sequence (see Fig. M4). A Fourier analysis of the oscillations reveals a dominant spectral peak at $f_{\nu}=2$. Taken together, these findings indicate that, within regime I, the integer and half-integer states belong to a single, unified quantum Hall sequence.
It is worth noting that the temperature dependence of the quantum Hall states closely mirrors that of the superconducting phase. As shown in Fig. 3d, the superconducting transition remains essentially unchanged across a broad range of $B_{\perp}$, and its transition temperature closely matches the onset temperature of the adjacent quantum Hall states.
Figure 4: Transition temperature and critical current. (a) Longitudinal resistance $R_{xx}$ (left panel) and transverse resistance $R_{xy}$ (right panel) measured as functions of carrier density $n$ and temperature $T$ across regime I at $B_{\perp}=3$ T and $D=1000$ mV/nm. (b) Differential resistance $dV/dI$ plotted as a function of $n$ and d.c. current bias $I_{\text{DC}}$ across regime I, measured at the same $B_{\perp}$ and $D$. Data in the left (right) panel were obtained in the longitudinal (transverse) measurement configuration.
Figure 4a presents an $n$$T$ map of the transport response across regime I, measured at fixed $B_{\perp}$ and displacement field $D$. The superconducting transition and the onset of the quantum Hall effect together trace a continuous envelope that evolves smoothly with carrier density $n$. A similar correspondence is observed in the current–voltage (I–V) characteristics. As shown in Fig. 4b, the superconducting state exhibits a peak in the differential resistance $dV/dI$, which defines a critical supercurrent that varies with $n$ across regime I.
Remarkably, the quantum Hall state on the high-density side also displays strong sensitivity to d.c. current bias: a dip in $dV/dI$ marks a critical current that separates a highly nonlinear response at low bias from a linear response at high bias. This critical current in the quantum Hall regime closely tracks the evolution of the critical supercurrent in the adjacent superconducting phase. Together, these observations strongly support a direct connection between the mechanisms underlying the superconducting and quantum Hall phases within regime I.
Together, our findings present a compelling challenge to existing models of superconductivity and quantum Hall physics. Any successful theoretical framework must account for the following key observations:
First, a superconducting phase is stabilized—rather than suppressed—by an out-of-plane magnetic field $B_{\perp}$. That the superconducting phase persists beyond $B_{\perp}>4$ T is unprecedented in two-dimensional superconductors. Second, a sequence of quantum Hall states emerges at both integer and half-integer filling factors, forming a single unified hierarchy. Third, the unique interplay between the superconducting and quantum Hall phases—including their shared onset temperature and adjacent locations in phase space—suggests a common underlying mechanism. Finally, the boundaries of regime I follow a Streda slope of $t=2.5$, consistent with a fractional quantum Hall state, yet without signatures of an incompressible state.
A possible interpretation of the Streda slope involves two components: a compressible state formed by carriers in a partially filled band, and a gapped, interaction-driven state characterized by a nontrivial Chern number of 2.52.52.5. In this picture, lines of constant Streda slope correspond to fixed charge density within the compressible band. The boundary of the cylindrical feature shown in Fig. 1b follows the same contour in the $D$–(charge density in the partially filled band) plane for each value of magnetic field.
Building on this framework, we now consider two candidate scenarios that may account for the observed superconducting and quantum Hall states.
The first scenario invokes the intriguing possibility of a quantum Hall effect of bosons. In this picture, the application of an out-of-plane magnetic field stabilizes pre-formed Cooper pairs—bound states of two electrons—within regime I. On the low-density side of regime I, a superconducting phase emerges as these bosonic pairs acquire long-range phase coherence. The remarkable stability of this superconducting phase against strong out-of-plane magnetic fields suggests that the Cooper pairs form in an odd orbital angular momentum channel, leading to an unconventional order parameter. On the high-density side, quantum Hall states arise when the bosonic Cooper pairs occupy their own Landau levels, in contrast to conventional models in which single electrons fill fermionic Landau levels. This bosonic quantum Hall scenario has been extensively explored in theoretical proposals  Cooper (2020) and realized in cold atom platforms  Léonard et al. (2023), but its manifestation in solid-state systems has remained elusive until now.
As shown in Fig. 3a–b, the boundary of the superconducting phase exhibits slight dips along the trajectories of the quantum Hall states, indicating that superconductivity is suppressed at the onset of quantum Hall states. This observation suggests that the superconducting and bosonic quantum Hall phases are competing orders within the same underlying bosonic framework.
This bosonic framework provides a natural explanation for the experimentally observed connection between superconducting and quantum Hall states: both emerge at similar magnetic fields, where Cooper pairs are stabilized, and both are suppressed at approximately the same temperature, $T\approx 0.5$ K, where Cooper pairs are destroyed. This also accounts for the $B$-independence of the onset temperature, as the Cooper pairing instability is not directly influenced by the magnetic field.
Moreover, the picture of competing bosonic phases offers a compelling explanation for the hierarchical structure of the quantum Hall sequence at both integer and half-integer filling factors—features that are less straightforward in a fermionic Landau level description. In fermionic systems, the energy scales for integer and fractional quantum Hall effects are governed by single-particle gaps and interaction energies, respectively, leading to marked differences in their robustness. In contrast, the quantum Hall effect in bosonic systems necessarily arises from interactions, implying that both integer and fractional bosonic states are governed by a common, interaction-driven energy scale. We note that filling factors discussed here are expressed in terms of electron density. Since Cooper pairs carry twice the charge of electrons, their effective filling factor is reduced by a factor of four.
In this picture, the Streda slope of $t=2.5$ observed in regime I reflects a topological gap located below the Fermi surface. While this topological structure is not directly responsible for the emergence of superconductivity and the quantum Hall effect, it modifies the charge carrier density and the dielectric environment in a manner that is favorable for stabilizing the Cooper pairing instability.
We also consider a second scenario in which a fractional Chern insulator (FCI) governs the physics of regime I. In this picture, regime I is populated by anyonic quasiparticles, as implied by the fractional Streda slope $t=2.5$. The quantum Hall states appearing at electronic filling $\nu=N/2$ is then interpreted as a sequence of quantum Hall states formed by these anyonic quasiparticles. In turn, superconductivity emerges as another correlated phase of anyons, leading to an anyonic superconductor—an exotic and theoretically rich phase that has recently attracted considerable interest Xu and et al. (2025); Nosov et al. (2025); Pichler et al. (2025). Within this framework, the characteristic onset temperature $T\approx 0.5~\textrm{K}$ reflects the energy scale necessary to stabilize the underlying FCI.
While this scenario appears consistent with most experimental observations, it faces several notable challenges. For example, effective Landau levels (LLs) of anyonic quasiparticles have not been identified to form such an extensive sequence, making their realization here highly nontrivial. Furthermore, filling LLs up to $\nu=6.5$ requires a substantial density of quasiparticles. Achieving such a high quasiparticle population while maintaining the robustness of a fractional Chern insulator (FCI) is highly unusual and presents a serious constraint for this scenario. In particular, this scenario raises a question regarding the Chern number of the underlying topological band.
Having considered two possible scenarios to account for our observations, we note that alternative explanations may still be viable. Nonetheless, the discovery of a superconducting phase that is stabilized—rather than suppressed—by an out-of-plane magnetic field, and its striking connection to an exotic sequence of quantum Hall states, represents a fundamentally new phenomenon. The intriguing interplay between superconductivity and quantum Hall effect presents a compelling challenge to existing theoretical paradigms, which is expected to stimulate broad future interest into emergent phenomena of strongly interacting electrons in 2D material heterostructures.
acknowledgments
This material is based on the work supported by the Air Force Office of Scientific Research under award no. FA9550-23-1-0482. R.Q.N., N.J.Z, and J.I.A.L. acknowledge support from the Air Force Office of Scientific Research. E.M. acknowledge support from U.S. National Science Foundation under Award DMR-2143384. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy. D.E.F. was supported in part by the National Science Foundation under grant No. DMR-2204635.