### v2

# A Hierarchy of Topological and Superconducting States in Rhombohedral Hexalayer Graphene

thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.
A Hierarchy of Topological and Superconducting States in Rhombohedral Hexalayer Graphene
Ron Q. Nguyen1    Hai-Tian Wu1    Erin Morissette1    Naiyuan J. Zhang1    Peiyu Qin1    K. Watanabe2    T. Taniguchi3    Aaron W. Hui1    D. E. Feldman1,5    J.I.A. Li1,4 jia.li@austin.utexas.edu 1Department of Physics, Brown University, Providence, RI 02912, USA 2Research Center for Functional Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 3International Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 4Department of Physics, University of Texas at Austin, Austin, TX 78712, USA 5Brown Theoretical Physics Center, Brown University, Providence, Rhode Island 02912, USA
(August 11, 2025)
Superconductivity and the quantum Hall effect are conventionally viewed as mutually exclusive: the former is suppressed by magnetic fields, while the latter relies on them. Here, we report the surprising coexistence of these two phenomena in rhombohedral hexalayer graphene. In this system, a superconducting phase is not destroyed—but instead stabilized—by an out-of-plane magnetic field. Strikingly, this superconducting state coexists and competes with a sequence of quantum Hall states that appear at both integer and half-integer Landau level fillings. Both the superconducting and quantum Hall states exhibit sharply defined thermal transitions or crossovers, with nearly identical onset temperatures—pointing to a shared underlying mechanism. Taken together, our observations uncover an unprecedented interplay between superconducting and topological phases, challenging conventional paradigms and opening a new frontier in condensed matter physics.
Figure 1: Multi-dimensional phase space of rhombohedral hexalayer graphene. (a) Schematic of the R6G heterostructure. The red arrow indicates the out-of-plane magnetic field $B_{\perp}$. (b) Schematic phase diagram defined by carrier density $n$, displacement field $D$, and out-of-plane magnetic field $B_{\perp}$. (c, d) nnDD maps of longitudinal resistance measured at $B=0$, with current flowing along the principal axes of (c) minimum and (d) maximum conductivity. (e) Annotated nnDD map highlighting prominent phases corresponding to panels (c, d). (f, g) nnBB maps measured at $D=967$ mV/nm, showing (f) longitudinal resistance $R_{xx}$ and (g) transverse resistance $R_{xy}$. White dashed lines trace the trajectories of integer quantum Hall states, exhibiting integer Streda slopes. The boundaries of regime I, highlighted by the black dashed lines, follow a Streda slope of $t=2.5$. In panel (f), white dotted lines show Streda slopes of 2 and 3, which clearly deviate from the boundaries of regime I.
Figure 2: $B_{\perp}$-stabilized superconductivity and coexisting quantum Hall state. nnDD maps of (a) $R_{xx}$ and (b) $R_{xy}$ measured at different values of $B_{\perp}$. The dashed line outlines the boundary of regime I, while the dash-dotted line encloses the superconducting phase. (c) Line cuts of $|R_{xy}|$ as a function of $n$ at $D=1000$ mV/nm, measured at various values of $B_{\perp}$. (d, e) Temperature dependence of $R_{xx}$ (red) and $R_{xy}$ (blue) measured from (d) the $B_{\perp}$-stabilized superconducting phase and (e) the adjacent quantum Hall state.
Figure 3: An exotic sequence of quantum Hall states. (a, b) nnBB map of $R_{xx}$ (top panel) and $R_{xy}$ (bottom panel) measured at (a) $D=980$ mV/nm and (b) $D=1000$ mV/nm. The superconducting region is enclosed by a white dashed line in the top panel, while the sequence of quantum Hall states forms a Landau fan, marked by black dashed lines.
An enduring theme in the study of superconductivity is its fundamental antagonism with magnetic fields: superconducting phases are typically suppressed by even modest applied fields. It is therefore striking when superconductivity persists under large magnetic fields, defying conventional wisdom. Such behavior sets high-$T_{c}$ superconductors apart from their more mundane counterparts, establishing a vibrant research frontier over the last decades  Phillips et al. (2021). Even more exotic is the scenario in which a superconducting phase is not merely resilient to magnetic fields, but is in fact stabilized by them—pointing to an unconventional pairing symmetry. A prominent example is UTe2, a candidate spin-triplet superconductor that exhibits field-induced reentrant superconductivity under strong magnetic fields Ran et al. (2019a, b); Aoki et al. (2019); Lewin et al. (2023).
When superconductivity persists to high magnetic fields, it opens the door to a rare and intriguing regime in which superconductivity coexists with another hallmark of two-dimensional (2D) electron systems: the quantum Hall effect. The interplay between these two phenomena has long been proposed as a key platform for realizing fault-tolerant quantum computation, owing to the potential emergence of non-Abelian anyons and topological superconductivity Nayak et al. (2008); Das Sarma et al. (2015); Peralta Gavensky et al. (2020); Alicea and Fendley (2016). Over the past decade, this coexistence has been explored extensively in engineered hybrid systems, where superconducting and quantum Hall regions are coupled across artificial interfaces Amet et al. (2016); Lee et al. (2017); Gül et al. (2022); Zhao et al. (2023), as well as in 2D material heterostructures hosting superconductivity and topological order at zero magnetic field Stepanov et al. (2020); Choi et al. (2025). Despite these advances, achieving an intrinsic intertwinement of superconductivity and the quantum Hall effect within a single 2D electronic system—without relying on interface engineering—remains a central open challenge. A recent experimental report has identified such coexistence near zero magnetic field, sparking renewed interest in the possibility of anyonic superconductivity and its topological implications  Xu et al. (2025); Nosov et al. (2025); Pichler et al. (2025); Shi and Senthil (2024).
Here, we report an unprecedented form of interplay between superconductivity and topological order under a large out-of-plane magnetic field, realized in a rhombohedral hexalayer graphene (R6G) heterostructure  Zhou et al. (2021a, b, 2022); Han et al. (2025); Choi et al. (2025). In stark contrast to all previously reported superconducting phases in 2D material heterostructures, the superconducting state in R6G is stabilized—rather than suppressed—by an external magnetic field. Strikingly, quantum Hall states emerge in direct proximity to this field-induced superconducting phase and form an unconventional sequence that includes both integer and half-integer Landau level (LL) fillings. Through multiple lines of experimental evidence, we identify a common microscopic origin underlying both phases. These findings place strong constraints on viable theoretical interpretations and point to a highly unconventional physical regime, where superconductivity and quantum Hall order emerge from a unified framework.
Facilitated by a dual-encapsulated heterostructure, as illustrated in Fig. 1a, the low-temperature phase space of R6G is governed by three independently tunable parameters: carrier density $n$, displacement field $D$, and out-of-plane magnetic field $B_{\perp}$. Superconductivity and the sequence of quantum Hall states are both confined within a unique boundary, which encircles a tilted cylindrical volume within this nnDDBB space, as illustrated in Fig. 1b. For simplicity, we will refer to this portion of the phase space as regime I.
We begin by defining the boundary of regime I through its distinct transport signatures. To this end, we perform measurements across multiple two-dimensional slices of the nnDDBB parameter space. In particular, we examine nnDD planes at five different values of out-of-plane magnetic field $B_{\perp}$ (blue dashed lines in Fig. 1b), as well as nnBB planes at three fixed values of displacement field $D$ (red dashed lines in Fig. 1b).
At zero magnetic field ($B_{\perp}=0$), the cross section of regime I is bounded by the electronic smectic phase and the striped superconducting state, both previously identified through angle-resolved transport studies Morissette et al. (2025a). The transport response across this regime is characterized in Figs. 1c–d, which show longitudinal resistance measured across the nnDD plane at $B_{\perp}=0$, along two orthogonal principal axes corresponding to maximum and minimum conductivity. The smectic phase occupies the region of pronounced transport anisotropy (see Fig. M2b), highlighted by the red-shaded area in Fig. 1e. Within this anisotropic regime, two superconducting phases—denoted SC i and SC ii—emerge concurrently with the smectic order and inherit its strong directional character  Morissette et al. (2025a).
The $B_{\perp}=0$ cross-section of regime I appears as a metallic phase situated between SC i and SC ii, enclosed by black dashed lines in Figs. 1c–e. As $B_{\perp}$ increases, the boundaries of regime I trace a well-defined trajectory in the nnBB plane, marked by peaks in the longitudinal resistance $R_{xx}$ and abrupt jumps in the Hall resistance $R_{xy}$. These boundaries, indicated by black dashed lines in Fig. 1f-g, exhibit a Streda slope of $t=2.5$  Středa and Smrčka (1983); Widom (1983); MacDonald and Středa (1984).
Within the stripe regime, two additional metallic phases—regime II and regime III—are highlighted in Figs. 1c–e. These phases resemble regime I in their transport characteristics and exhibit well-defined boundaries in the nnBB plane, following distinct Streda slopes: $t=2$ for regime II and $t=3.5$ for regime III (see Fig. M8 for further discussion). The presence of such sharply delineated, evolving boundaries is intriguing. Although a well-defined Streda slope often signals a topological phase, we do not observe transport features typically associated with incompressible quantum Hall states Huang et al. (2025). We revisit this issue later in the context of additional experimental evidence.
In addition to their distinct Streda slopes, regimes I, II, and III exhibit significantly reduced transport anisotropy compared to the surrounding stripe phase (see Fig. M2c and Fig. M3a), suggesting the emergence of a competing order that disrupts the underlying smectic state.
Figures 2a–b show nnDD maps of the longitudinal and Hall resistances measured at various values of $B_{\perp}$. At each field, regime I is delineated by peaks in $R_{xx}$, defining a boundary that shifts systematically with $B_{\perp}$ in both carrier density $n$ and displacement field $D$. These evolving boundaries are marked by dashed lines in Figs. 2a–b. Most notably, upon applying a non-zero $B_{\perp}$, a superconducting pocket emerges on the low-density side of regime I, while a sequence of quantum Hall states appears on the high-density side.
The superconducting phase—enclosed by dash-dotted lines—is identified by vanishing longitudinal resistance $R_{xx}$ and zero Hall resistance $R_{xy}$, appearing as dark blue and white, respectively, in the color maps. Figure 2c presents line cuts of the transport response as a function of $n$, taken at fixed $D=1000$ mV/nm and varying $B_{\perp}$. At $B_{\perp}=0$, the system remains resistive across the entire density range. In contrast, at $B_{\perp}=2.5$ and $3.3$ T, a superconducting dome is clearly visible, characterized by a sharp drop to zero resistance. Upon further increasing $B_{\perp}$ to $5$ T, the superconducting phase is fully suppressed.
On the high-density side of the superconducting region, we observe the emergence of a distinct phase along the upper boundary of regime I, exhibiting transport signatures that sharply contrast with those of the superconducting state. In this regime, vanishing $R_{xx}$ is accompanied by a pronounced peak in $R_{xy}$. The corresponding conductivity matrix is characterized by vanishing diagonal (longitudinal) components and finite, antisymmetric off-diagonal (Hall) components. In Fig. 3, we show that this high-density state evolves with varying $B_{\perp}$ along a well-defined Streda slope—a hallmark of an incompressible quantum Hall state. Notably, the minima in $R_{xx}$ for each quantum Hall state are flanked by resistive peaks, suggesting that these states arise from strong electronic interactions.
Remarkably, both the superconducting and quantum Hall phases exhibit sharply defined onsets as a function of temperature (Fig. 2d–e). The RRTT curves in Fig. 2d display behavior consistent with a superconducting transition, characterized by a critical temperature $T_{c}\approx 0.5$ K. In contrast, the similarly sharp onset observed for the quantum Hall state in Fig. 2e is highly unusual. Conventional quantum Hall states typically display a gradual thermal activation in $R_{xx}$, while $R_{xy}$ remains largely quantized and temperature-independent (see Fig. M6). The abrupt temperature dependence observed here suggests an unconventional origin for the quantum Hall state, a possibility we return to in light of additional experimental evidence discussed below.
Next, we investigate the unusual robustness of the superconducting phase against $B_{\perp}$ in Figs. 3a–b, which present transport measurements across nnBB planes intersecting regime I at fixed displacement fields. A pronounced resistive peak in $R_{xx}$ appears within regime I around $B_{\perp}=1$ T, marking the lower-field boundary for both the superconducting and quantum Hall states. We interpret this feature as indicative of a transition in the underlying electronic order.
Given that the superconducting phase is stabilized above this resistive feature, it is natural to consider a transition in the underlying isospin order. For $B_{\perp}>1$ T, regime I is expected to be fully polarized in both spin and valley degrees of freedom. In this high-field regime, superconductivity likely originates from Cooper pairing in an odd orbital angular momentum channel, consistent with a $p$-wave pairing symmetry Anderson and Morel (1961). At lower fields ($B_{\perp}<1$ T), the ground state may be only partially polarized. Fermiology analysis reveals that regime I lies between a half-metallic phase at low density and a quarter-metallic phase at high density, consistent with partial isospin polarization in the intermediate regime. However, this interpretation may be at odds with the chiral nature of the SC i and SC ii phases observed near zero field Han et al. (2025), which have been argued to emerge from a fully isospin-polarized normal state.
An alternative interpretation is that the resistive feature marks a $B_{\perp}$-driven reconstruction of momentum-space order. In rhombohedral graphene (R6G), rotational symmetry breaking has been attributed to spontaneous polarization in momentum space Morissette et al. (2025b, a); Dong et al. (2023); Jung et al. (2015); Huang et al. (2023). The stability of the superconducting phase may therefore depend sensitively on the momentum-space configuration of the Fermi surface, particularly within a quarter-metallic state  Geier et al. (2024); Parra-Martinez et al. (2025). This possibility is further supported by the observed $B_{\perp}$-dependent rotation of the principal axis within regime I (see Fig. M3).
On the high-density side of regime I, quantum Hall states emerge along well-defined linear trajectories in the $n\text{--}B$ plane, tracing minima in $R_{xx}$ and maxima in $|R_{xy}|$. These trajectories, marked by black dashed lines in Fig. 3a–b, form a characteristic pattern reminiscent of a Landau fan diagram Zhang et al. (2005). However, the sequence of quantum Hall states in this regime exhibits several unusual features that set it apart from conventional Landau fans.
In particular, the observed trajectories follow Streda slopes corresponding to both integer and half-integer Landau level (LL) fillings, with $\nu=N/2$, where $N\in\mathbb{Z}$. This pattern would conventionally imply the emergence of incompressible states at fully and half-filled LLs. However, a striking discrepancy arises between the inferred filling factors and the actual Hall quantization. As shown in Fig. 3c–e, line cuts along representative trajectories reveal vanishing $R_{xx}$, signaling fully developed incompressible states, while the Hall resistance $R_{xy}$ remains quantized at integer multiples of the conductance quantum.
This mismatch between the Streda slope and Hall quantization suggests an unconventional origin. One possibility is a reduction in the effective LL degeneracy by a factor of two, such that a sequence of integer quantum Hall states occurs along Streda slope at $\nu=N/2$.
In addition, the observed quantum Hall states exhibit a highly unusual hierarchical structure. In conventional two-dimensional electron systems, incompressible states at half-filled Landau levels—commonly associated with even-denominator fractional quantum Hall effects (FQHE)—possess significantly smaller energy gaps than their integer counterparts and exhibit strong dependence on the out-of-plane magnetic field $B_{\perp}$. By contrast, quantum Hall states in regime I display comparable robustness between integer and half-integer Streda slopes, evidenced by nearly identical onset temperature, as illustrated in Fig. 4a. This offers further indication that the underlying incompressible states are IQHE in nature.
Moreover, the onset temperature of quantum Hall states remain nearly constant across a wide range of $B_{\perp}$. This temperature dependence closely mirrors that of the superconducting phase. As shown in Fig. 4b, the superconducting transition remains essentially unchanged across a broad range of $B_{\perp}$, and its transition temperature closely matches the onset temperature of the adjacent quantum Hall states.
Figure 4: Transition temperature and critical current. (a, b) Temperature dependence of $R_{xy}$ measured at different $B_{\perp}$ values within (a) quantum Hall states and (b) the superconducting phase. (c) Longitudinal resistance $R_{xx}$ (left panel) and transverse resistance $R_{xy}$ (right panel) as functions of carrier density $n$ and temperature $T$ across the superconducting and quantum Hall states at $B_{\perp}=3$ T and $D=1000$ mV/nm. (d) Differential resistance $dV/dI$ as a function of $n$ and d.c. current bias $I_{\mathrm{DC}}$ at $B_{\perp}=3$ T and $D=1000$ mV/nm. Data in the left (right) panel were obtained in the longitudinal (transverse) measurement configuration. (e–g) Schematic diagrams illustrating possible mechanisms underlying superconductivity and quantum Hall states. (e) Regime I is associated with the emergence of a CDW order, with insulating islands shown as white circles. (f, g) Away from the low-density boundary of regime I, itinerant electrons occupy the regions between insulating islands, giving rise to (f) superconductivity on the low-density side of regime I and (g) the quantum Hall effect on the high-density side of regime I.
A close correlation between the quantum Hall and superconducting states is further illustrated in Fig. 4c–d. An $n\text{--}T$ map of the transport response, taken at fixed $B_{\perp}$ and displacement field $D$, reveals that the superconducting transition and the onset of the quantum Hall effect trace a continuous envelope that evolves smoothly with carrier density $n$. A similar correspondence is observed in the current–voltage (I–V) characteristics, as shown in Fig. 4d. The quantum Hall state, which appears on the high-density side of regime I, exhibits pronounced nonlinearity in its transport response, evidenced by large variations in the differential resistance $dV/dI$ in both longitudinal and transverse channels. This nonlinearity defines a sharp threshold in the d.c. current bias, which closely aligns with the critical current of the superconducting state—marked by a peak in $dV/dI$. Taken together, these coordinated behaviors strongly suggest a shared origin underlying both the superconducting and quantum Hall states.
In the following, we discuss potential mechanisms that may underlie the intriguing coexistence of superconductivity and quantum Hall states. Any viable explanation must account for the following key experimental observations. First, a superconducting phase is stabilized—rather than suppressed—by an out-of-plane magnetic field $B_{\perp}$. That the superconducting phase persists beyond $B_{\perp}>4$ T is unprecedented in two-dimensional superconductors. Second, the sequence of quantum Hall states emerges at both integer and half-integer filling factors, displaying a prominent mismatch between the Streda slope and Hall quantization. Third, the unique interplay between the superconducting and quantum Hall phases—including their shared onset temperature and adjacent locations in phase space—suggests a common underlying mechanism.
We consider a scenario involving charge density wave (CDW) order, wherein electrons are localized to form an array of insulating islands, schematically illustrated as white circles in Fig. 4e–g. In this picture, the extent of regime I is governed by the stability of the CDW order across the $n\text{--}D\text{--}B$ phase space. As the carrier density $n$ increases from the low-density boundary of regime I, itinerant electrons are introduced occupying the regions between the insulating islands. These itinerant electrons contribute to conductive transport at $B_{\perp}=0$ and enable the emergence of superconductivity and quantum Hall states at $B_{\perp}>1$ T. The boundary of the CDW phase traces a finite Streda slope in the $n\text{--}B$ plane, suggesting that the reconstructed electronic bands feature non-zero Berry curvature. However, due to the mismatch between Streda slope and Hall plateau, the exact Streda slope likely is not representing the actual Berry curvature.
In the presence of an out-of-plane magnetic field, LL formation of the itinerant electrons gives rise to the sequence of quantum Hall states observed in regime I (Fig. 4f). Within this framework, these quantum Hall states resemble a re-entrant integer quantum Hall effect Eisenstein et al. (2002); Xia et al. (2004); Halperin and Jain (2020). Because electrons localized within the insulating islands do not contribute to transport, the effective carrier density is reduced. This reduction yields Hall plateaus comparable to those of the IQHE emanating from the CNP, indicating the same effective density occurring at distinct values of the actual carrier density (Fig. 2b and Fig. 3a–b).
Crucially, the emergence of CDW order decouples the Streda slope from the quantized Hall plateau. When extra charge carriers are added to the sample, they preferentially occupy the regions between insulating islands. As a result, the effective density of itinerant electrons, $n^{\ast}$, increases at a faster rate compared to the actual electron density $n$. This naturally accounts for the mismatch between the Streda slope and the quantized Hall plateau.
This framework leads to an intriguing scenario in which superconductivity coexists with charge density wave (CDW) order, occupying the regions between insulating CDW islands (Fig. 4g). These metallic regions supports superconducting transport at low temperatures. The close proximity of the superconducting and quantum Hall phases in the phase diagram—and their nearly identical onset temperatures—suggests that both originate from the same population of itinerant electrons, with superconductivity (quantum Hall states) stabilized in the low-density (high-density) limit. As the temperature increases, the CDW order melts, disrupting the spatial segregation between localized and itinerant carriers. This melting transition naturally accounts for the sharply defined thermal onset observed in both superconducting and quantum Hall responses (Fig. 4a–b).
Besides the CDW scenario, we note that alternative explanations may still be viable. Nonetheless, the discovery of a superconducting phase that is stabilized—rather than suppressed—by an out-of-plane magnetic field, and its striking connection to an exotic sequence of quantum Hall states, represents a fundamentally new phenomenon. The intriguing interplay between superconductivity and quantum Hall effect presents a compelling challenge to existing theoretical paradigms, which is expected to stimulate broad future interest into emergent phenomena of strongly interacting electrons in 2D material heterostructures.
acknowledgments
This material is based on the work supported by the Air Force Office of Scientific Research under award no. FA9550-23-1-0482. R.Q.N., N.J.Z, and J.I.A.L. acknowledge support from the Air Force Office of Scientific Research. E.M. acknowledge support from U.S. National Science Foundation under Award DMR-2143384. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy. D.E.F. was supported in part by the National Science Foundation under grant No. DMR-2204635.